Given the following information about your firm’s capital structure, calculate your firm’s WACC (assume the corporate tax rate is 35%).

  • Debt
    • Number of bonds outstanding = 14,000
    • price per bond = $1,165
    • par value per bond = $1,000
    • coupon rate = 6% (paid annually)
    • Years to maturity = 10

  • Common Stock
    • Number of shares outstanding = 1,000,000
    • Price per share = $25
    • Book value per share = $15
    • Beta = 1.4
    • Risk free rate = 4.5%
    • Market risk premium = 5%

A.) 7.19%

B.) 8.53%

C.) 7.72%

D.) 7.98%

E.) 9.08%
Answer #1

Coupon = 0.06 * 1000 = 60

YTM = 3.9689%

Keys to use in a financial calculator: FV 1000, PV -1165, N 10, PMT 60, CPT I/Y

Market value of bonds = 14,000 * 1165 = 16,310,000

Cost of equity = risk free rate + beta (market risk premium)

Cost of equity = 0.045 + 1.4 (.05)

Cost of equity = 0.115 or 11.50%

Market value of common stock = 1,000,000 * 25 = 25,000,000

Total market value = 16,310,000 + 25,000,000 = 41,310,000

WACC = (25,000,000 / 41,310,000)*0.115 + (16,310,000 / 41,310,000)*0.039689*(1 - 0.35)

WACC = 0.069596 + 0.01019

WACC = 0.0798 or 7.98%